# Mechanically concealed holes

### v2

@@ V Conclusions
Indeed, a corresponding expression for the thickness of the shell around the hole was identified and derived as a function of the mechanical material parameters and the radius of the concealed hole. For this purpose, we assumed continuous, linearly elastic, homogeneous, isotropic materials under plane-strain conditions. The concentrations of the holes were low enough so that mechanical interactions between the holes can be neglected. One step further, we extended this possibility of mechanical concealment from a macroscopic continuum perspective down to discrete atomistic scales using molecular dynamics simulations. A purely atomistic, planar model solid was considered. Although gradual deviations in our finite-sized atomistic systems emerge at lower shell stiffnesses, which necessitate thicker shells when particulate inhomogeneities play a larger role, overall consistent results were found. In principle, they validate the concept introduced via continuum elasticity theory down to the atomistic level.
@@
Future extensions of this work are manifold, important, and obvious. An upcoming investigation shall address more general types of loading, beyond uniform. Besides plane-strain, also plane-stress and three-dimensional situations shall be considered in future works. Elevated concentrations of holes that make them mutually interact through their induced deformations shall be evaluated. More complex considerations include mechanically anisotropic materials. Altogether, we wish to support the further development of quantitative measures in the context of light-weight materials design.
@@ Acknowledgements.
The generated molecular dynamics simulation data, used to derive the results presented herein, including all the figures, are openly available in the repository Zenodo and can be found at https://doi.org/10.5281/zenodo.17957498.
@@ Appendix A Lennard-Jones “toy model” for a strongly cohesive solid with a hole
Choosing specific combinations of materials in atomistic simulations in our specific geometry of a hole, shielded by an elastic shell, embedded in a surrounding elastic solid, limits, for each realization, the comparison between continuum theory and atomistic simulations to a specific ratio of the material parameters $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$. To this end, we choose to model a physically traceable system that is simple enough to retain isotropy (potential without angular dependence), yet complex enough to allow mechanical concealment using a shell of variable stiffness around a hole.
@@
Consistently with the chosen value of $\sigma$, all atomistic systems simulated in this work possess a constant lattice parameter $r_{0}=2^{1/6}\sigma$. This can be inferred from the radial distribution function g(r) computed for different systems with and without holes, see Fig. 12.

### v3

@@ IV Conclusions
Indeed, a corresponding expression for the thickness of the shell around the hole was identified and derived as a function of the mechanical material parameters and the radius of the concealed hole. For this purpose, we assumed continuous, linearly elastic, homogeneous, isotropic materials under plane-strain conditions. The concentrations of the holes were low enough so that mechanical interactions between the holes can be neglected. One step further, we extended this possibility of mechanical concealment from a macroscopic continuum perspective down to discrete atomistic scales using molecular dynamics simulations. A purely atomistic, planar model solid was considered. Overall, consistent results were found. They validate the concept introduced via continuum elasticity theory down to the atomistic level.
@@
Future extensions of this work are manifold, important, and obvious. First, we remark that our results within continuum elasticity theory directly carry over to uniform expansion of the system, which, however, may be less abundant in practical situations. An upcoming investigation shall address more general types of loading, beyond uniform. Specifically, our next steps concern uniaxial strains and shear. To lowest order, this extension will allow us to include mutual interactions between shielded holes. Besides plane-strain, also plane-stress and three-dimensional situations shall be considered in future works. Elevated concentrations of holes that make them mutually interact through their induced deformations shall be evaluated. More complex considerations include mechanically anisotropic materials. Altogether, we wish to support the further development of quantitative measures in the context of light-weight materials design.
@@ Acknowledgements.
The generated molecular dynamics simulation data, used to derive the results presented herein, including the figures, are openly available in the repository Zenodo and can be found at https://doi.org/10.5281/zenodo.18470981.
@@ Appendix A Lennard-Jones “toy model” for a strongly cohesive solid with a hole
Choosing specific combinations of materials in atomistic simulations in our specific geometry of a hole, shielded by an elastic shell, embedded in a surrounding elastic solid, limits, for each realization, the comparison between continuum theory and atomistic simulations to a specific ratio of the material parameters $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$.
To this end, we choose to model a physically traceable system that is simple enough to retain isotropy (potential without angular dependence), yet complex enough to allow mechanical concealment using a shell of variable stiffness around a hole.
@@
[…]